Use the following information to calculate the requested Income Statement metrics.

Net Farm Income

$36,000

Opportunity Cost of Labor

$16,000

Average Equity

$220,000

Opportunity Cost of Management

$8,000

Average Asset Value

$360,000

Family Living Expenses

$20,000

Interest Expense

$11,000

Income and SS taxes

$4,000

Total Revenue

$109,500

Opportunity Cost of Capital

10%

a. Rate of Return on Assets

b. Rate of Return on Equity

c. Operating Profit Margin Ratio

d. Change in Equity

e. Return to Management
Answer #1

a. Rate of return on assets= 6.38%

Rate of return on assets (%)= retrun to assets x 100 average assets

Need to calculate return on Asset

  Return to assets = (Net farm income + interest expenses) less: (opprotunity cost of labor + opportunity cost of management)

  Return to assets = S(36,000 + 11,000 - $(16,000+8,000) = $47,000 - $24,000 = $23,000

Average assets value = $360,000. (as per the question)

Rate of return on assets (%)= $23,000 -x100 $360,000 = 6.388% = 6.39% (rounded off)

b.Rate of return on equity __5.45 %

Rate of return on equity (%) retrun to equity -x100 average equity

Need to calculate return on equity

Return on equity Net farm income less: (opprotunity cost of labor + opportunity cost of management)

Return on equity = $36,000 - $(16,000 +8,000) =$36,000-$24,000 =$12,000

Average equity = $220,000. (as per the question).

Rate of return on equity (%)= $12,000 x 100 $220,000 = 5.454% = 5.45% (rounded off)

c.  Operating profit margin ratio ____21 %

Operating profit margin ratio operating profit gross / total revenue

need to calculate operating profit

   Operating profit = (Net farm income + interest expenses) less: (opprotunity cost of labor + opportunity cost of management)

Operating profit = $(36,000+ 11,000) -S(16,000+8,000) = $47,000 - $24,000 = $23,000

total revenue = $109,500

     $23,000 Operating profit margin ratio = $109,500 = 0.2100 = 21% (rounded off)

d.  Change in equity $ __$ 12,000_______

Net farm income less : Family living expenses income and social security taxes = Change in market-basis equity

Change in Equity = $36,000-$20,000-$4,000

=$12,000

e.  Return to management $ ____$6,000

Return on management = Return to labor and management less : opprotunity cost of labor

Need to calculate return on labour & management:

  Net farm income from operations less : opprotunity cost of equity capital

Return to labor and management = $36,000 - (10%$220,000) = $14,000

  Return on management = $14,000 - $8,000 =$6,000